# Basic representations of genus zero nonabelian Hodge spaces

### v2

@@ 2.1.3. Modified formal data
2.2. Admissible deformations, fission data, and forests
We now recall a few definitions and results about admissible deformations of irregular connections, involving the notions of levels and fission trees, and diagrams. We refer the reader to [11, 18] for more details.
2.2.1. Local version
@@
A full irregular type at $a\in\Sigma$ of rank $n$ is a Galois-closed list $T=(q_{1},\dots,q_{n})$ of exponential factors in some direction $d\in\partial$ (i.e. $q_{i}$ is a section of $\mathcal{I}_{a}$ in a germ of sector around $d$).
Any full irregular type $T=(q_{1},\dots,q_{n})$ defines an associated irregular class $\Theta$ in the following way: its active circles are the connected components $\langle q_{i}\rangle$ in $\mathcal{I}$ of the exponential factors, and the multiplicity $n_{i}$ of $\langle q_{i}\rangle$ is the multiplicity of $q_{i}$ in the list $(q_{1},\dots,q_{n})$. Passing from a full irregular type to an irregular class amounts to forgetting the ordering of the exponential factors.
@@
Let $T=(q_{1},\dots,q_{n})$ and $T^{\prime}=(q^{\prime}_{1},\dots,q^{\prime}_{n})$ be rank $n$ full irregular types at $a$ in some direction $d$. We say that $T$ and $T^{\prime}$ are admissible deformations of each other, and denote this by $T\sim T^{\prime}$, if we have
$\operatorname{slope}(q_{i}-q_{j})=\operatorname{slope}(q^{\prime}_{i}-q^{\prime}_{j}),$
for all $i,j=1,\dots,n$.
@@
In particular, if $\Theta\sim\Theta^{\prime}$, denoting by $N$ and $N^{\prime}$ the sets of active circles of $\Theta$, $\Theta^{\prime}$ respectively, there exists a bijection $\phi:N\to N^{\prime}$ such that for each active circle $I\in N$, $\phi(I)$ is an admissible deformation of $I$, and $I$, $I^{\prime}$ have the same multiplicity. We say that such a bijection is compatible.
@@
Let $(\Theta,\mathcal{C})$ and $(\Theta^{\prime},\mathcal{C}^{\prime})$ two sets of local formal data. We say that they are admissible deformations of each other if $\Theta\sim\Theta^{\prime}$ and there exists a compatible bijection $\phi:N\to N^{\prime}$ between their respective sets of active circles $N,N^{\prime}$ such that $\mathcal{C}^{\prime}_{\phi(I)}=\mathcal{C}_{I}$ for all $I\in N$.
@@
(2.1) $\mathrm{Levels}(\Theta)=\{\operatorname{slope}(q_{\alpha}-q_{\beta})\;|\;\alpha,\beta\in\mathtt{I}_{d}\}\setminus\{0\}\subset\mathbb{Q}_{>0}$
@@
Let us first consider the case of an irregular class having only one Stokes circle. Let $I$ be a Stokes circle at $a$. We can always write $I=\langle q\rangle_{a}$, with $q=\sum_{i=1}^{p}a_{i}z_{a}^{-k_{i}}$, for some integer $p$, with $k_{1}>\dots>k_{p}$ and $a_{i}\neq 0$ for every $i=1,\dots,p$. The rational numbers $k_{i}$ are uniquely defined by $I$, and we call them the exponents of $I$. We have the following:
Proposition 2.6.
a) Two Stokes circles $I,J\subset\mathcal{I}_{a}$ are admissible deformations of each other if and only if $\operatorname{Levels}(I)=\operatorname{Levels}(J)\subset\mathbb{Q}$.
b) A subset $(k_{1}>k_{2}>\cdots>k_{m})\subset\mathbb{Q}_{>0}$ is the set of levels of some circle $I\subset\mathcal{I}_{a}$ if and only if
(2.2) $k_{1},k_{2},\ldots,k_{m}$ have strictly increasing common denominators $>1$.
In other words if $d_{i}$ is the denominator of $k_{i}$ (in lowest terms) and
(2.3) $r_{i}\text{ is the lowest common multiple of }d_{1},d_{2},\ldots,d_{i}$
for each $i$ (so that $r_{i}\;|\;r_{i+1}$), then $1<r_{1}<r_{2}<\cdots<r_{m}$.
c) Let $I$ be a Stokes circle, and let $(k_{1}>\dots>k_{p})$ denote its set of exponents. Then $\operatorname{Levels}(I)$ is the largest possible subset of $\{k_{1},\dots,k_{p}\}$ of the form $k_{i_{1}}>\dots>k_{i_{m}}$ for with $1=i_{1}<\dots<i_{m}\leq p$ such that $k_{i_{1}},\dots,k_{i_{m}}$ have strictly increasing common denominators $>1$.
Passing to the general case with irregular classes having several Stokes circles, to characterise admissible deformations, we need more data than the levels of each individual Stokes circle. To this end we introduce the notion of common part and fission exponent of a pair of Stokes circles.
@@
In turn, if $I=\langle q\rangle$ is a Stokes circle, we set $I_{\geq k}:=\langle q_{\geq k}\rangle$. We define in a similar way truncations $I_{>k}$, $I_{\leq k}$, $I_{<k}$.
@@
The fission exponent of $I$ and $I^{\prime}$ is the rational number
$f_{I,\widehat{I}}=\max(\operatorname{slope}(I_{<k}),\operatorname{slope}(\widehat{I}_{<k}))\in\frac{1}{r\widehat{r}}\mathbb{Q}.$
If $I=\widehat{I}$ then $I_{c}=I=\widehat{I}$ and $f_{I,\widehat{I}}=0$, otherwise $f_{I,\widehat{I}}>0$. $I_{c}=0$ if and only if the leading terms of $I$ and $\widehat{I}$ have different Galois orbits, in this case we say that $I$ and $\widehat{I}$ have no common part.
Lemma 2.8.
Let $I$ and $\widehat{I}$ be two Stokes circles with ramification orders $r,\widehat{r}$. The set of slopes among the rational numbers $\operatorname{slope}({q_{i}}-{\widehat{q}_{j}})$, for $i=0,\dots,r-1$, $j=0,\dots,\widehat{r}-1$ is of the form
$\operatorname{Levels}(I_{c})\sqcup\{f_{I,\widehat{I}}\}\subset\mathbb{Q}_{\geq 0},$
i.e. it consists of the levels of the common part of together with their fission exponent.
This motivates the following definition:
@@
A (local) “fission datum” is a pair $\mathcal{F}=(\mathcal{L},f)$ where $\mathcal{L}$ is a multiset
$\mathcal{L}=L_{1}+\cdots+L_{n}$
of level data and $f$, the fission exponents, are the choice of a rational number $f_{ij}=f_{ji}\in\mathbb{Q}_{\geq 0}$, for all $i,j\in\{1,\ldots,n\}$.
A (local) irregular class determines a fission datum as follows. If $\Theta=\sum_{1}^{n}I_{i}$ is a rank $n$ irregular class at some point $a$ (where the Stokes circles $I_{i}$ are not necessarily distinct), then define $L_{i}=L(I_{i})$ to be the level datum of $I_{i}$ for each $i$, and define $\mathcal{L}(\Theta):=\sum_{1}^{n}L_{i}$ to be the corresponding multiset of level data. Then by taking $f_{ij}=f_{I_{i},I_{j}}$ to be corresponding fission exponents, this determines the fission datum $\mathcal{F}(\Theta)=(\mathcal{L}(\Theta),f)$ of the irregular class $\Theta$. Note that the multiplicity of any given Stokes circle $I_{j}$ in the class $\Theta$ is determined from the fission data by the recipe:
$\Theta(I_{j})=\absolutevalue{\{i=1,2,\ldots,n\;|\;f_{ij}=0\}}.$ | .
@@
Theorem 2.10.
Two (local) irregular classes $\Theta$ and $\Theta^{\prime}$ are admissible deformations of each other if and only if they have the same fission datum.
Notice that not any fission datum $\mathcal{F}$ actually comes from an irregular class $\Theta$. To characterise such fission data, as well as provide a more graphical way of encoding them, the notion of a (twisted) fission tree is introduced in [11]. We will not be needing to recall the exact axioms for fission trees here, so we refer the reader once more to loc. cit. for the full details. Let us just briefly summarize the construction of the fission tree $\mathcal{T}(\Theta)$ associated to an irregular class $\Theta$.
@@
    To each active circle $I$ of $\Theta$ with level datum $L$ one associates a full branch $T_{I}$, which is the datum of a copy of $\mathbb{R}_{\geq 0}$, and of a set of distinguished points on it, the vertices, defined as the set of all possible exponents of a Stokes circle with level datum $L$. The real coordinate on a full branch will be called the height.
@@
    The tree $\mathcal{T}(\Theta)$ is the topological space obtained by gluing for each pair of disctinct active circles $I,J$ the corresponding full branches together above a height defined as the smallest height of any vertex in some full branch strictly greater than the fission exponent $f_{I,J}$.
In brief, the leaves of the tree are in bijection with the active circles of $\Theta$, and the tree contains the information about the level data as well as the fission exponents. In particular the datum of the fission tree $\mathcal{T}(\Theta)$ is equivalent to the datum of $\mathcal{F}(\Theta)$.
Finally, if we include conjugacy classes of formal monodromies, a set of local formal data $(\Theta,\mathcal{C})$ determines a fission tree $\mathbf{T}:=\mathcal{T}(\Theta)$, and $\mathcal{C}$ amounts to the datum of a conjugacy class for each leaf of $\mathbf{T}$. Class of admissible deformations of local formal data is characterised by such pairs $(\mathbf{T},\mathcal{C})$.
2.2.2. Global version
Until now in this paragraph we have dealt only with the local situation. The global case is described in a similar way, by collecting together the local data at each singularity.
@@ 2.2.2. Global version
Corollary 2.11.
Two global irregular classes $\mathbb{P}^{1}$ are admissible deformations of each other if and only if they have the same global fission datum, or equivalently the same fission forest.
As a consequence, a class of admissible deformations of wild Riemann surfaces with boundary data is characterised by a pair $(\mathbf{F},\bm{\mathcal{C}})$, where $\mathbf{F}$ is a fission forest, and $\mathcal{C}$ the datum of a conjugacy class for each leaf of $\mathbf{F}$.
2.3. Diagrams
Another invariant that we can attach to any connection $(E,\nabla)$ on a Zariski open subset of $\mathbb{P}^{1}$ is a diagram $\Gamma(E,\nabla)$ [18]. As for the forest, the diagram only depends on the (modified) formal data $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ of the connection.
The diagram $\Gamma(E,\nabla)=\Gamma(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})=(N,B)$ of the connection $(E,\nabla)$ has the following structure: it consists of a core diagram $\Gamma_{c}(E,\nabla)=\Gamma_{c}(\bm{\breve{\Theta}})$ with nodes $N$, to which is then glued, onto each core node, a single leg determined by a conjugacy class.
The set of core nodes $N$ is the set of Stokes circles of the modified irregular class $\bm{\breve{\Theta}}$. In terms of the non-modified irregular class $\bm{\Theta}$, $N$ is the disjoint union of three finite sets:
$N=$$\{\text{Stokes circles of $(E,\nabla)$ at $\infty$}\}\ \cup$
$\{\text{wild Stokes circles of }(E,\nabla)\text{ not at $\infty$}\}\ \cup$
{tame Stokes circles of ($\{\text{tame Stokes circles of }(E,\nabla)\text{ not at $\infty$ with nontrivial formal monodromy}\}.$
The matrix of edge multiplicities between the core nodes is then defined as follows. For any Stokes circle $I=\langle q\rangle$, let $\alpha_{I}=\text{Irr}(I),\beta_{I}=\operatorname{ram}(I)\in\mathbb{N}$ denote its irregularity and ramification numbers, so that $I$ has slope $\alpha_{I}/\beta_{I}$. We define an integer $B^{\infty}_{IJ}$ as follows:
(2.4) $B^{\infty}_{IJ}$ $=A_{IJ}-\beta_{I}\beta_{J},\qquad\text{if $I\neq J$,}$
(2.5) $B^{\infty}_{II}$ $=A_{II}-\beta_{I}^{2}+1,\quad\,\text{otherwise,}$,
where $A_{IJ}=\operatorname{Irr}(\operatorname{Hom}(I,J))$ is the irregularity of the irregular class $\operatorname{Hom}(I,J)$.

### v3

@@ 2.1.3 Modified formal data
2.2 Admissible deformations, fission data, and forests
We now recall a few definitions and results about admissible deformations of irregular connections, involving the notions of levels and fission trees, and about diagrams. We refer the reader to [13, 20] for more details and for proofs of the statements.
2.2.1 Local version
@@
A full irregular type at $a\in\Sigma$ of rank $n$ is a Galois-closed list $T=(q_{1},\dots,q_{n})$ of exponential factors in some direction $d\in\partial$ (i.e., $q_{i}$ is a section of $\mathcal{I}_{a}$ in a germ of sector around $d$).
Any full irregular type $T=(q_{1},\dots,q_{n})$ defines an associated irregular class $\Theta$ in the following way: its active circles are the connected components $\bigl\langle q_{i}\bigr\rangle$ in $\mathcal{I}$ of the exponential factors, and the multiplicity $n_{i}$ of $\bigl\langle q_{i}\bigr\rangle$ is the multiplicity of $q_{i}$ in the list $(q_{1},\dots,q_{n})$. Passing from a full irregular type to an irregular class amounts to forgetting the ordering of the exponential factors.
@@
Let $T=(q_{1},\dots,q_{n})$ and $T^{\prime}=(q^{\prime}_{1},\dots,q^{\prime}_{n})$ be rank $n$ full irregular types at $a$ in some direction $d$. We say that $T$ and $T^{\prime}$ are admissible deformations of each other, and denote this by $T\sim T^{\prime}$, if we have $\operatorname{slope}(q_{i}-q_{j})=\operatorname{slope}(q^{\prime}_{i}-q^{\prime}_{j})$, for all $i,j\in\{1,\dots,n\}$.
@@
In particular, if $\Theta\sim\Theta^{\prime}$, denoting by $N$ and $N^{\prime}$ the sets of active circles of $\Theta$ and $\Theta^{\prime}$, respectively, there exists a bijection $\phi\colon N\to N^{\prime}$ such that for each active circle $I\in N$, $\phi(I)$ is an admissible deformation of $I$, and $I$, $I^{\prime}$ have the same multiplicity. We say that such a bijection is compatible.
@@
Let $(\Theta,\mathcal{C})$ and $(\Theta^{\prime},\mathcal{C}^{\prime})$ be two sets of local formal data. We say that they are admissible deformations of each other if $\Theta\sim\Theta^{\prime}$ and there exists a compatible bijection $\phi\colon N\to N^{\prime}$} between their respective sets of active circles $N$, $N^{\prime}$ such that $\mathcal{C}^{\prime}_{\phi(I)}=\mathcal{C}_{I}$ for all $I\in N$}.
@@
$\mathrm{Levels}(\Theta)=\{\operatorname{slope}(q_{\alpha}-q_{\beta})\mid\alpha,\beta\in\mathtt{I}_{d}\}\setminus\{0\}\subset\mathbb{Q}_{>0},$
@@
Let us first consider the case of an irregular class having only one Stokes circle. Let $I$ be a Stokes circle at $a$. We can always write $I=\bigl\langle q\bigr\rangle_{a}$, with $q=\sum_{i=1}^{p}a_{i}z_{a}^{-k_{i}}$, for some integer $p$, with $k_{1}>\dots>k_{p}$} and $a_{i}\neq 0$ for every $i\in\{1,\dots,p\}$. The rational numbers $k_{i}$ are uniquely defined by $I$, and we call them the exponents of $I$. We have the following.
Proposition 2.7 ([13, Proposition 3.1]).
  • $(a)$
    Two Stokes circles $I,J\subset\mathcal{I}_{a}$ are admissible deformations of each other if and only if $\operatorname{Levels}(I)=\operatorname{Levels}(J)\subset\mathbb{Q}$.
  • $(b)$
    A subset $(k_{1}>k_{2}>\cdots>k_{m})\subset\mathbb{Q}_{>0}$ is the set of levels of some circle $I\subset\mathcal{I}_{a}$ if and only if $k_{1},k_{2},\ldots,k_{m}$ have strictly increasing common denominators >1>1. In other words if $d_{i}$ is the denominator of $k_{i}$ (in lowest terms)) and $r_{i}$ is the lowest common multiple of $d_{1},d_{2},\ldots,d_{i}$} for each $i$ (so that $r_{i}\mid r_{i+1})$, then $1<r_{1}<r_{2}<\cdots<r_{m}$.
  • $(c)$
    Let $I$ be a Stokes circle, and let $(k_{1}>\dots>k_{p})$ denote its set of exponents. Then $\operatorname{Levels}(I)$ is the largest possible subset of $\{k_{1},\dots,k_{p}\}$ of the form $k_{i_{1}}>\dots>k_{i_{m}}$ for with $1=i_{1}<\dots<i_{m}\leq p$ such that $k_{i_{1}},\dots,k_{i_{m}}$ have strictly increasing common denominators >1>1.
Passing to the general case with irregular classes having several Stokes circles, to characterise admissible deformations, we need more data than the levels of each individual Stokes circle. To this end, we introduce the notion of common part and fission exponent of a pair of Stokes circles.
@@
In turn, if $I=\bigl\langle q\bigr\rangle$ is a Stokes circle, we set $I_{\geq k}:=\bigl\langle q_{\geq k}\bigr\rangle$ (this is well-defined since it does not depend on the choice of representative). We define in a similar way truncations $I_{>k}$, $I_{\leq k}$, $I_{<k}$.
@@
The fission exponent of $I$ and $\widehat{I}$ is the rational number
$f_{I,\widehat{I}}=\max\bigl(\operatorname{slope}(I_{<k}),\operatorname{slope}\bigl(\widehat{I}_{<k}\bigr)\bigr)\in\frac{1}{r\widehat{r}}\mathbb{Q}$.
If $I=\widehat{I}$, then $I_{c}=I=\widehat{I}$ and $f_{I,\widehat{I}}=0$, otherwise $f_{I,\widehat{I}}>0$. $I_{c}=0$ if and only if the leading terms of $I$ and $\widehat{I}$ have different Galois orbits, in this case we say that $I$ and $\widehat{I}$ have no common part.
Lemma 2.9 ([13, Lemma 3.11]).
Let $I$ and $\widehat{I}$ be two Stokes circles with ramification orders $r^\widehat{r}$. The set of slopes among the rational numbers $\operatorname{slope}\bigl({q_{i}}-{\widehat{q}_{j}}\bigr)$, for $i\in\{0,\dots,r-1\}$, $j\in\big\{0,\dots,\allowbreak{\widehat{r}-1}\big\}$ is of the form $\operatorname{Levels}(I_{c})\sqcup\{f_{I,\widehat{I}}\}\subset\mathbb{Q}_{\geq 0}$, i.e., it consists of the levels of the common part of the two Stokes circles together with their fission exponent.
This motivates the following definition.
@@
A (local) fission datum is a pair $\mathcal{F}=(\mathcal{L},f)$ where $\mathcal{L}$ is a multiset $\mathcal{L}=L_{1}+\cdots+L_{n}$ of level data and $f$, the fission exponents, consist of the choice of a rational number $f_{ij}=f_{ji}\in\mathbb{Q}_{\geq 0}$, for all $i,j\in\{1,\ldots,n\}$.
A (local) irregular class determines a fission datum as follows. If $\Theta=\sum_{1}^{n}I_{i}$ is a rank $n$ irregular class at some point $a$ (where the Stokes circles $I_{i}$ are not necessarily distinct), then define $L_{i}=\mathrm{Levels}(I_{i})$ to be the level datum of $I_{i}$ for each $i$, and define $\mathcal{L}(\Theta):=\sum_{1}^{n}L_{i}$ to be the corresponding multiset of level data. Then by taking $f_{ij}=f_{I_{i},I_{j}}$ to be corresponding fission exponents, this determines the fission datum $\mathcal{F}(\Theta)=(\mathcal{L}(\Theta),f)$ of the irregular class $\Theta$. Note that the multiplicity of any given Stokes circle $I_{j}$ in the class $\Theta$ is determined from the fission data by the recipe $|\Theta(I_{j})=\absolutevalue{\{i\in\{1,2,\ldots,n\}\mid f_{ij}=0\}}$.
@@
Theorem 2.11 ([13, Theorem 3.17]).
Two ((local)) irregular classes $\Theta$ and $\Theta^{\prime}$ are admissible deformations of each other if and only if they have the same fission datum.
Notice that not any fission datum $\mathcal{F}$ actually comes from an irregular class $\Theta$. To characterise such fission data, as well as provide a more graphical way of encoding them, the notion of (twisted) fission tree was introduced in [13]. We will not be needing to repeat all the exact axioms for fission trees here, so we just briefly recall the main properties of the fission tree associated to an irregular class, referring the reader once more to loc. cit. for full details, and to Section 4 for many concrete examples of fission trees.
If $\Theta$ is an irregular class at infinity, its fission tree $\mathcal{T}(\Theta)$ is a tuple $(\mathcal{T},\mathbb{V},\mathbb{A},\mathbb{L},h,n)$, where
@@
    $\mathcal{T}$ is a metrized tree, with set of vertices $\mathbb{V}$. An edge of the tree is an element of $E:=\pi_{0}(\mathcal{T}\smallsetminus\mathbb{V})$ of $\mathcal{T}$. Any vertex that is not a leaf nor the root is adjacent to $\geq 2$ edges, one of which is the parent edge, and the others are the descendant edges. The branch vertices $\mathbb{Y}\subset\mathbb{V}$} are those with $\geq 2$ descendants. The trunk of the tree is the union of all edges and vertices above all the branch vertices.
    If $i$ is a leaf of $\mathcal{T}$, we define the full branch $\mathcal{B}_{i}$ of $i$ as the subset of $\mathcal{T}$ corresponding to the path from $i$ all the way to the end of the trunk.
@@
    $\mathbb{L}$ and $\mathbb{A}$ are two subsets of $\mathbb{V}$, whose elements are respectively called mandatory and admissible vertices, and $\mathbb{L}\subset\mathbb{A}$. The set $\mathbb{L}$ is finite and can be empty.
  • $\bullet$
    $h\colon\mathcal{T}\!\to\!\mathbb{R}_{\geq 0}$ is a length-preserving function, which we call the height, such that $h^{-1}(0)\!=:\!\mathbb{V}_{0}$} is the set of leaves of $\mathcal{T}$, $h$ maps isomorphically any edge to an interval, and for any leaf $i$, $h$ maps the full branch $\mathcal{B}_{i}$ isomorphically onto $\mathbb{R}_{\geq 0}$.
  • $\bullet$
    $n$ is a map $\mathbb{V}_{0}\to\mathbb{Z}_{\geq 1}$, i.e., consists of the datum of an integer multiplicity $n_{i}\geq 1$ for each leaf $i\in\mathbb{V}_{0}$.
Moreover, the fission tree $\mathcal{T}(\Theta)$ has the following properties:
  • $\bullet$
    The leaves are in one-to-one correspondence with the Stokes circles of $\Theta$: if $\Theta=\sum_{i\in N}n_{i}I_{i}$, we have $\mathbb{V}_{0}=N$, with the leaf $i$ corresponding to the Stokes circle $I_{i}$, and having multiplicity $n_{i}$.
  • $\bullet$
    If $i\in\mathbb{V}_{0}$ is a leaf and $I$ is the corresponding Stokes circle, the heights of the mandatory vertices of the full branch $\mathcal{B}_{i}$ are exactly the levels of $I$, that is $h(\mathbb{L}\cap\mathcal{B}_{i})=\operatorname{Levels}(I)$.
  • $\bullet$
    If $i,j\in\mathbb{V}_{0}$ are two distinct leaves, $I$, $J$ are the corresponding Stokes circles, and $v_{ij}\in\mathbb{V}$ is closest common ancestor of $i$, $j$ then the height of all children of $v_{ij}$ is equal to the fission exponent $f_{I,J}$.
In brief, the leaves of the tree correspond to the Stokes circles of $\Theta$, and the tree contains the information about the level data as well as the fission exponents. In particular, the datum of the fission tree $\mathcal{T}(\Theta)$ is equivalent to the datum of $\mathcal{F}(\Theta)$.
Finally, if we include conjugacy classes of formal monodromies, a set of local formal data $(\Theta,\mathcal{C})$ determines a fission tree $\mathbf{T}:=\mathcal{T}(\Theta)$, and $\mathcal{C}$ amounts to the datum of a conjugacy class in $\mathrm{GL}_{n}(\mathbb{C})$ for each leaf of $\mathbf{T}$, where $n$ is its multiplicity. In turn, a class of admissible deformations of local formal data is characterised by such pairs $(\mathbf{T},\mathcal{C})$.
2.2.2 Global version
Until now, in this paragraph we have dealt only with the local situation. The global case is described in a similar way, by collecting together the local data at each singularity.
@@ 2.2.2 Global version
Corollary 2.12 ([13, Corollary 3.33]).
Two global irregular classes on $\Sigma$ are admissible deformations of each other if and only if they have the same global fission datum, or equivalently the same fission forest.
As a consequence, a class of admissible deformations of wild Riemann surfaces with boundary data is characterised by a pair $(\mathbf{F},\bm{\mathcal{C}})$, where $\mathbf{F}$ is a fission forest, and $\bm{\mathcal{C}}$ the datum of a conjugacy class in $\operatorname{GL}_{n}(\mathbb{C})$ for each leaf of $\mathbf{F}$, where $n$ is the leaf multiplicity.
2.3 Diagrams
Another invariant that we can attach to any connection $(E,\nabla)$ on a Zariski open subset of $\mathbb{P}^{1}$ is a diagram $\Gamma(E,\nabla)$ [20]. As for the forest, the diagram only depends on the (modified) formal data $\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$ of the connection.
The diagram $\Gamma(E,\nabla)=\Gamma\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)=(N,B)$ of the connection $(E,\nabla)$ has the following structure: it consists of a core diagram $\Gamma_{c}(E,\nabla)=\Gamma_{c}\bigl(\bm{\breve{\Theta}}\bigr)$ with nodes $N$, to which is then glued, onto each core node, a leg determined by a conjugacy class.
The set of core nodes $N$ is the set of Stokes circles of the modified irregular class $\bm{\breve{\Theta}}$. In terms of the non-modified irregular class $\bm{\Theta}$, $N$ is the disjoint union of three finite sets
[…]
The matrix of edge/loop multiplicities between the core nodes is then defined as follows. For any Stokes circle $I=\bigl\langle q\bigr\rangle$, let $\alpha_{I}=\text{Irr}(I),\beta_{I}=\operatorname{ram}(I)\in\mathbb{N}$ denote its irregularity and ramification numbers, so that $I$ has slope $\alpha_{I}/\beta_{I}$. We define an integer $B^{\infty}_{IJ}$ as follows:
$B^{\infty}_{IJ}:=A_{IJ}-\beta_{I}\beta_{J}\quad\text{if}\ I\neq J,\qquad B^{\infty}_{II}:=A_{II}-\beta_{I}^{2}+1\quad\text{otherwise,}$
where $A_{IJ}:=\operatorname{irr}(\operatorname{Hom}(I,J))$ is the irregularity of the irregular class $\operatorname{Hom}(I,J)$.